OpenStudy (cheesecakekitten):

Which expression is equivalent to 4^-2?

OpenStudy (cheesecakekitten):

will give medal and fan

OpenStudy (anonymous):

\(\bbox[8pt,#f9bec7,border:5pt solid #33cccc]{\color{white}{\Large\cal~\bigstar\heartsuit\dagger What~can~I~help~you~with~today?\bigstar\heartsuit\dagger{{\cal\color{#33cccc}{\hspace-281pt{\bigstar\heartsuit\dagger What~can~I~help~you~with~today?\bigstar\heartsuit\dagger}}}}~~}}\) \(\small{Made~by~K~Lynn}\)

OpenStudy (cheesecakekitten):

A. 1/4×1/4 B. (4)(−2) C. 1/2×1/2×1/2×1/2 D.(−2)(−2)(−2)(−2)

OpenStudy (cheesecakekitten):

OOPS!

OpenStudy (cheesecakekitten):

hold on a sec

OpenStudy (cheesecakekitten):

A. 1/4 x 1/4 B. (4) (-2) C. 1/2 x 1/2 x 1/2 x 1/2 D. (-2)(-2)(-2)(-2)

OpenStudy (cheesecakekitten):

there

OpenStudy (displayerror):

\[4^{-2} = \left( \frac{1}{4} \right)^2\] So which choice represents this?

OpenStudy (cheesecakekitten):

A?

OpenStudy (cheesecakekitten):

?

OpenStudy (cheesecakekitten):

hello?

OpenStudy (displayerror):

Yep! When you have a number raised to a negative exponent, just flip the number and make the exponent positive. Here are some examples: \[\left( \frac{1}{5} \right)^{-3} = 5^3\] \[\left( \frac{a}{b} \right)^{-c} = \left( \frac{b}{a} \right)^c\]

OpenStudy (anonymous):

Correct! :D

OpenStudy (cheesecakekitten):

yay!

OpenStudy (cheesecakekitten):

thnx

OpenStudy (cheesecakekitten):

baibai